Have I simplified the below correctly?
Problem 1:
3x^2 + 6x - 45 = x^2 + 2x - 15 = (x - 3) (x + 5)= answer of x-3 after cancellation
3x^2 + 21x +30 X^2 + 7x + 10 (x + 2) (x + 5) x-2
Problem 2:
4a^2b^4 / 8ab^2 = 4A^2b = 4a^2b^4 * 1 = 4a^2b^4 = 4ab^2 = simplified to 1ab^2
3c^2 3c^2 8ab^2 24ab^2c^2 24c^2 6c^2
Problem 3:
16x^2 - 1 / 4x + 1 = 16x^2 * y^3 = simplified to 4xy
y^2 y^3 y^2 4x + 1
Thank you

Found 2 solutions by greenestamps, MathTherapy:
Answer by greenestamps(13203)



Again it is impossible to figure out what the original problems are, or what you are doing with them....


Answer by MathTherapy(10556)

Have I simplified the below correctly?
Problem 1:
3x^2 + 6x - 45 = x^2 + 2x - 15 = (x - 3) (x + 5)= answer of x-3 after cancellation
3x^2 + 21x +30 X^2 + 7x + 10 (x + 2) (x + 5) x-2
Problem 2:
4a^2b^4 / 8ab^2 = 4A^2b = 4a^2b^4 * 1 = 4a^2b^4 = 4ab^2 = simplified to 1ab^2
3c^2 3c^2 8ab^2 24ab^2c^2 24c^2 6c^2
Problem 3:
16x^2 - 1 / 4x + 1 = 16x^2 * y^3 = simplified to 4xy
y^2 y^3 y^2 4x + 1
Thank you
You're NOT presenting your expressions in a way that they can be easily understood.
Anyway, based on the last one, I think you mean the following:
Problem # 1:    <====== This is what you have.

                <====== This is what you should have.

Compare them and you'll see where you went wrong!!


Problem # 2:    <====== I THINK this is what you have.

                <====== This is what you should have, if my HUNCH is right.

Compare them and you'll see where you went wrong!!


Problem # 3:   <====== I THINK this is what you have.

               <==== This is what you should have, if my HUNCH is right.

Compare them and you'll see where you went wrong!!